OpenStudy (anonymous):

please help! Use the substitution method to solve the following system of equations. 4x – y – 3z = –3 x – 2y – 2z = –3 x + y + 3z = –2 (–1, 2, –1) (1, –2, 1) (–1, –2, 1) (1, –2, –1)

OpenStudy (anonymous):

\[\begin{cases}4x-y-3z=-3\\x-2y-2z=-3\\ x+y+3z=-2\end{cases}\] Solve for one variable in terms of the other two: \[\begin{cases}x=\frac{y+3z-3}{4}\\\\x=2y+2z-3\\\\ x=-y-3z-2\end{cases}\] First two equations give you \[\frac{y+3z-3}{4}=2y+2z-3\\ y+3z-3=8y+8z-12\\ -7y-5z=-9\\ \color{red}{7y+5z=9}\] and the last two equations give you \[2y+2z-3=-y-3z-2\\ \color{red}{3y+5z=1}\] Solve the red system, then from there solve for x.

OpenStudy (anonymous):

how do i solve the red system?

OpenStudy (anonymous):

Subtract 3y+5z=1 from 7y+5z=9 . The result is 4y = 8.

OpenStudy (anonymous):

Or you can keep up with the substitution method: \[\begin{cases}7y+5z=9\\ 3y+5z=1\end{cases}\] Solving the second equation for \(5z\), you have \[5z=1-3y\] Substituting into the first equation, you get \[7y+(1-3y)=9\\ 4y=8\] Either way works

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

what is z i have only heard of x,y